how can adding doubles help you multiply by 2?
Answer #1
Adding doubles is the same thing as multiplying by 2. Instead of memorizing the two tables, all you have to do is add the same number twice. Examples: 4*2=4+4=8, 7*2=7+7=14, etc.
answered by: cain